OpenStudy (anonymous):

Write an equation in standard form of the line that passes through (0,3) (2,-3)

OpenStudy (ranga):

From the given two points, find the slope m Put one of the points and the m value in y = mx + b and find b. Rearrange the line equation so it looks like: Ax + By = C which will be the standard form.

OpenStudy (anonymous):

so would it be 3x+y=5 or no?

OpenStudy (ranga):

no. if you put either of the two points (0,3) or (2,-3) into your equation it does not satisfy either one. so that is not the correct equation.

OpenStudy (ranga):

3x + y = 3